Calculate the simple interest rate at which one can earn $1,500 per day interest on an investment of $7,500,000
Answer by Alan3354(69443)

Calculate the simple interest rate at which one can earn $1,500 per day interest on an investment of $7,500,000
-----------
1500/7500000 = 1.5e3/7.5e6 = 0.2e-3 = 2e-4
% = 2e-4*100 = 2e-2% per day
= 0.02% per day
0.02*365 = 7.3% per year